### v1

# Runtime Analysis of the SMS-EMOA for Many-Objective Optimization

Abstract
The widely used multiobjective optimizer NSGA-II was recently proven to have considerable difficulties in many-objective optimization. In contrast, experimental results in the literature show a good performance of the SMS-EMOA, which can be seen as a steady-state NSGA-II that uses the hypervolume contribution instead of the crowding distance as the second selection criterion.
This paper conducts the first rigorous runtime analysis of the SMS-EMOA for many-objective optimization. To this aim, we first propose a many-objective counterpart, the $m$-objective $m$OJZJ problem, of the bi-objective OJZJ benchmark, which is the first many-objective multimodal benchmark used in a mathematical runtime analysis. We prove that SMS-EMOA computes the full Pareto front of this benchmark in an expected number of $O(M^{2}n^{k})$ iterations, where $n$ denotes the problem size (length of the bit-string representation), $k$ the gap size (a difficulty parameter of the problem), and $M=(2n/m-2k+3)^{m/2}$ the size of the Pareto front. This result together with the existing negative result on the original NSGA-II shows that in principle, the general approach of the NSGA-II is suitable for many-objective optimization, but the crowding distance as tie-breaker has deficiencies.
We obtain three additional insights on the SMS-EMOA. Different from a recent result for the bi-objective OJZJ benchmark, the stochastic population update often does not help for $m$OJZJ. It results in a $1/\Theta(\min\{Mk^{1/2}/2^{k/2},1\})$ speed-up, which is $\Theta(1)$ for large $m$ such as $m>k$. On the positive side, we prove that heavy-tailed mutation still results in a speed-up of order $k^{0.5+k-\beta}$. Finally, we conduct the first runtime analyses of the SMS-EMOA on the bi-objective OneMinMax and LOTZ benchmarks and show that it has a performance comparable to the GSEMO and the NSGA-II.
1 Introduction
The NSGA-II is the most widely-applied multiobjective evolutionary algorithm (MOEA). Non-dominated sorting and crowding distance are its two major features differentiating it from basic MOEAs such as the GSEMO or the $(\mu+1)$ SIBEA. Zheng, Liu, and Doerr [ZLD22] conducted the first runtime analysis of the NSGA-II (see [ZD23a] for the journal version). This work quickly inspired many interesting follow-up results in bi-objective optimization [ZD22, BQ22, DQ23a, DQ23b, DQ23c, DOSS23a, DOSS23b, ${}^{+}$23]. In contrast to these positive results for two objectives, Zheng and Doerr [ZD23b] proved that for $m\geq 3$ objectives the NSGA-II needs at least exponential time (in expectation and with high probability) to cover the full Pareto front of the $m$-objective OneMinMax benchmark, a simple many-objective version of the basic OneMax problem where all search points are Pareto optimal. They claimed that the main reason for this low efficiency is the independent computation of the crowding distance in each objective.
A very recent work showed that the NSGA-III, a successor algorithm of the NSGA-II aimed at better coping with many objectives, can efficiently solve the $3$-objective OneMinMax problem [WD23]. Since apparently practitioners much prefer the NSGA-II (more than 17,000 citations on Google scholar only in the last five years) over the NSGA-III (5,035 citations since its publication in 2013), it remains an interesting question whether there are variants of the NSGA-II which better cope with many objectives.
With the SMS-EMOA, an interesting variant of the NSGA-II was proposed by eume, Naujoks, and Emmerich [BNE07]. This algorithm is a steady-state variant of the NSGA-II (that is, in each iteration only a single offspring is generated and possibly integrated into the population) that further replaces the crowding distance as secondary selection criterion with the classic hypervolume contribution. Many empirical works (see the almost 2,000 papers citing [BNE07]) confirmed the good performance of the SMS-EMOA for many-objective optimization. The only mathematical runtime analysis of the SMS-EMOA was conducted only very recently by Bian et al. [BZLQ23]. They proved that its runtime for the bi-objective OJZJ problem is $O(n^{k+1})$ and proposed a stochastic population update mechanism which was proven to achieve an expected runtime of $O(n^{k+1}\min\{1,n/2^{k/4}\})$.
Our Contributions: This paper conducts the first mathematical runtime analysis of the SMS-EMOA for more than two objectives. We first define the $m$OJZJ benchmark, an $m$-objective counterpart of the bi-objective OJZJ problem [DZ21], which is the problem analyzed in the only previous runtime analysis for the SMS-EMOA [BZLQ23]. We note that the $m$OJZJ problem is the first multimodal many-objective benchmark proposed for a theoretical analysis, to the best of our knowledge. We prove that the SMS-EMOA covers the full Pareto front of this benchmark in an expected number of $O(M^{2}n^{k})$ iterations, where $n$ is the problems size, $k$ the gap size (a difficulty parameter of the problem), $m$ the number of objectives, and $M=(2n/m-2k+3)^{m/2}$ the size of the Pareto front.
We recall that the original NSGA-II needs at least exponential time to optimize the $m$OneMinMax problem, which is a special case of $m$OJZJ with gap size $k=1$. Since the SMS-EMOA employs non-dominated sorting, but replaces the crowding distance in the original NSGA-II by the hypervolume contribution, our result in a similar fashion as the analysis of the NSGA-III in [WD23] suggests that the general approach of the NSGA-II is suitable for many-objective optimization and that it is only the crowding distance as tie-breaker which is not appropriate for more than two objectives.
We then analyze whether the better performance of the SMS-EMOA on the bi-objective OJZJ problem achieved via a new stochastic population update [BZLQ23] extends to the $m$-objective $m$OJZJ problem. Unfortunately, we shall observe that only a speed-up of order $1/\Theta(\min\{Mk^{1/2}/2^{k/2},1\})$ is obtained, which is $\Theta(1)$ when $m$ is large, e.g., $m>k$.
On the positive side, we show that the advantage of heavy-tailed mutation is preserved. We analyze the SMS-EMOA with heavy-tailed mutation on the $m$OJZJ benchmark and prove that a speed-up of order $k^{0.5+k-\beta}$ is achieved. This is the same speed-up as observed for single-objective and bi-objective Jump problems [DLMN17, DZ21, ZD23c]. We note that this is the first theoretical work to support the usefulness of heavy-tailed mutation in many-objective optimization.
Finally, since so far the performance of the SMS-EMOA was only analyzed on the bi-objective OJZJ problem [BZLQ23], we conduct mathematical runtime analyses of the SMS-EMOA also on the two most prominent bi-objective benchmarks OneMinMax and LOTZ. We prove that the SMS-EMOA finds the Pareto fronts of these benchmarks in an expected number of iterations of at most $2e(n+1)n(\ln n+1)$ for OneMinMax and at most $2en^{2}(n+1)$ for LOTZ. These are the same asymptotic runtimes (in terms of fitness evaluations) as known for the GSEMO and the NSGA-II.
This paper is organized as follows. Section 2 briefly introduces some basic concepts for the many-objective optimization and the mathematical notations that will be used in this paper. The many-objective OJZJ benchmark is proposed in Section 3. Sections 4 to 6 analyze the runtime of the original SMS-EMOA, the SMS-EMOA with stochastic population update, and the one with heavy-tailed mutation on this many-objective benchmark. The runtime for two well-known bi-objective functions is conducted in Section 7. Section 8 concludes this work.
2 Preliminaries
In many-objective optimization, one tries to find good solutions for a problem containing several, but at least three objectives. This work considers pseudo-Boolean maximization, hence our problem is described by a function $f=(f_{1},\dots,f_{m}):\{0,1\}^{n}\rightarrow\mathbb{R}^{m}$. Here $m>2$ is the number of objectives and $n\in\mathbb{N}$ is called the problem size. As in bi-objective optimization (that is, the case $m=2$), usually not all solutions are comparable. We say that a solution $x\in\{0,1\}^{n}$ dominates a solution $y$, denoted as $x\succ y$, if $f_{i}(x)\geq f_{i}(y)$ for all $i\in\{1,\dots,m\}$, and at least one of these inequalities is strict. For a given many-objective function $f$, we say that $x\in\{0,1\}^{n}$ is a Pareto optimum if and only if there is no $y\in\{0,1\}^{n}$ that dominates $x$ w.r.t. $f$. The set of all Pareto optima is called Pareto set, and the set of all function values of the Pareto optima is called Pareto front. The goal for an MOEA is to find the full Pareto front as much as possible, that is, to compute a not too large set $S$ of solutions such that $f(S)$ equals or approximates well the Pareto front. As common in the mathematical runtime analysis of MOEAs, we call the runtime of an algorithm the number of function evaluations until its population $P$ covers the full Pareto front, that is, $f(P)$ contains the Pareto front. We refer to [NW10, AD11, Jan13, ZYQ19, DN20] for general introductions to the mathematical runtime analysis of evolutionary algorithms, and to [Bro11] for a discussion of runtime analyses of MOEAs.
Mathematical Notations. For $a,b\in\mathbb{N}$ with $a\leq b$, we denote the set of integers in the interval $[a,b]$ by $[a..b]=\{a,a+1,\dots,b\}$. For a point $x\in\{0,1\}^{n}$, let $x_{[a..b]}:=(x_{a},x_{a+1},\dots,x_{b})$ and $|x|_{1}$ be the number of ones in $x$.
3 The $m$-Objective Jump Function
As mentioned earlier, the NSGA-II was proven to have enormous difficulties in optimizing many-objective problems [ZD23b]. In that paper, the $m$-objective counterpart $m$OneMinMax of the bi-objective OneMinMax benchmark was proposed and analyzed. Since the only runtime work of the SMS-EMOA so far [BZLQ23] analyzed the bi-objective jump problem OJZJ, whose special case with gap size $k=1$ is (essentially) the OneMinMax problem, we shall propose and work with an $m$-objective version of OJZJ as well. Again, its special case $k=1$ will be (essentially) equal to the $m$OneMinMax problem. With this, our results are comparable both the ones in [ZD23b] and [BZLQ23].
3.1 $m$OJZJ
We first recall the definition of the $m$OneMinMax problem. For the ease of presentation, we only consider even numbers $m$ of objectives here. In the $m$OneMinMax problem, the bit string (of length $n$) is divided into $m/2$ blocks of equal length $2n/m$. On each of these, a bi-objective OneMinMax problem is defined. We note that this general block construction goes back to the seminal paper of [LTZ04].
Definition 1 ([ZD23b]).
Let $m$ be the number of objectives and be even, and the problem size $n$ be a multiple of $m/2$. Let $n^{\prime}=2n/m\in\mathbb{N}$. For any $x=(x_{1},\dots,x_{n})$, the $m$-objective function mOneMinMax $f:\{0,1\}^{n}\rightarrow\mathbb{R}^{m}$ is defined by
$f_{i}(x)$ =$=\begin{cases}\textsc{OneMax}(\bar{x}_{[\frac{i-1}{2}n^{\prime}+1..\frac{i+1}{2}n^{\prime}]}),&\text{if $i$ is odd}\\ \textsc{OneMax}({x}_{[\frac{i-2}{2}n^{\prime}+1..\frac{i}{2}n^{\prime}]}),&\text{else,}\end{cases}$
where $\bar{x}=(1-x_{1},\dots,1-x_{n})$ and the function $\textsc{OneMax}:\{0,1\}^{n^{\prime}}\rightarrow\mathbb{R}$ is defined by
$\textsc{OneMax}(y)=\sum_{i=1}^{n^{\prime}}y_{i}$
for any $y\in\{0,1\}^{n^{\prime}}$.
In order to include $m$OneMinMax as a special case of $m$OJZJ, we define the $m$-objective OJZJ in a similar manner, that is, we divide the $n$ bit positions into $m/2$ blocks and define a OJZJ problem in each block.
Definition 2.
[…]
$f_{i}(x)$ =$=\begin{cases}\textsc{Jump}_{n^{\prime},k}(x_{[\frac{i-1}{2}n^{\prime}+1..\frac{i+1}{2}n^{\prime}]}),&\text{if $i$ is odd}\\ \textsc{Jump}_{n^{\prime},k}(\bar{x}_{[\frac{i-2}{2}n^{\prime}+1..\frac{i}{2}n^{\prime}]}),&\text{else,}\end{cases}$
where $\bar{x}=(1-x_{1},\dots,1-x_{n})$ and the function $\textsc{Jump}_{n^{\prime},k}:\{0,1\}^{n^{\prime}}\rightarrow\mathbb{R}$ is defined by
[…]
for any $y\in\{0,1\}^{n^{\prime}}$.
We note that the function $\textsc{Jump}_{n^{\prime},k}$ used in the definition above is the famous Jump benchmark. It was first defined by Droste, Jansen, and Wegener [DJW02] and has quickly become the most employed multimodal benchmark in the theory of randomized search heuristics, leading to many fundamental results on how these algorithms cope with local optima [JW02, ${}^{+}$18, Doe21, BBD21, HS22, RW22, Wit23, DEJK24].
[…]
3.2 Characteristics
We now give more details on this $m$OJZJ function. Let $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}]$, $i\in[1..m/2]$, be the $i$-th block in the partition of $n$ bit positions. From Definition 2, we know that the bit values in the block $B_{i}$ only influence the objectives $f_{2i-1}$ and $f_{2i}$. Figure 1 plots the objective values of $f_{2i-1}$ and $f_{2i}$ relative to the number of ones in this block. Obviously, $m$OJZJ is multimodal with respect to the definition of multimodality of multiobjective problems in [DZ21].
Figure 1: The objective values of $f_{2i-1}$ and $f_{2i}$ in $m$OJZJ w.r.t. $|x_{B_{i}}|_{1}$, the number of ones in the block $B_{i}$.
It is not difficult to see that the Pareto set is $S^{*}=\{x\in\{0,1\}^{n}\mid|x_{B_{i}}|_{1}\in[k..n-k]\cup\{0,n\},i\in[1..m/2]\}$. Hence, the Pareto front is
[…]
${}\mid a_{1},\dots,a_{m/2}\in[2k..n^{\prime}]\cup\{k,n^{\prime}+k\}\}$
and the Pareto front size is $M:=|F^{*}|=(n^{\prime}-2k+3)^{m/2}$.
We recall at this point that we defined the runtime of an MOEA as the time to cover the full Pareto front $F^{*}$, that is, the first time to reach a population $P$ with $F^{*}\subseteq f(P)$.
Note that for $k>1$, not all points are Pareto optimal and not all sets of mutually non-dominated points are the subsets of the Pareto front. However, we have the following result showing that the maximum number of mutually non-dominated points is at most the size of the Pareto front.
Lemma 3.
Let $k\leq n^{\prime}/2$ and $P$ be a set of solutions such that any two are not weakly dominating (w.r.t. $m$OJZJ). Then $|P|\leq M$.
Proof.
Note that all objective values of any non-Pareto optimal solution are less than $2k$. Since each Pareto optimum has its all objective values of at least $2k$, we know that any non-Pareto optimal solution is dominated by all Pareto optima excluding the ones with some block all zeros or all ones (called extreme-block Pareto optima). Hence, $P$ excluding extreme-block Pareto optima, denoted as $P^{\prime}$, cannot include both Pareto optima and non-Pareto optimal solutions. If $P$ includes only Pareto optima, then we know that $|P|\leq M$.
We now consider that $P^{\prime}$ includes only non-Pareto optimal solutions, that is, for any $x\in P^{\prime}$, there exists a block $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$ such that $|x_{B_{i}}|_{1}\in[1..k-1]\cup[n^{\prime}-k+1..n^{\prime}-1]$. Note that for any block $B_{i}$, any two solutions $x,y$ are comparable w.r.t. $(f_{2i-1},f_{2i})$ if $(|x_{B_{i}}|_{1}\in[1..k-1],|y_{B_{i}}|_{1}\in[1..k-1])$ or $(|x_{B_{i}}|_{1}\in[n^{\prime}-k+1..n^{\prime}-1],|y_{B_{i}}|_{1}\in[n^{\prime}-k+1..n^{\prime}-1])$. Hence, with the bit values on other blocks fixed, there are at most two solutions in $P^{\prime}$ that are not weakly dominated. We also note that $(f_{2i-1}(z),f_{2i}(z))>(f_{2i-1}(x),f_{2i}(x))$ holds if $(z_{B_{i}}=1^{n^{\prime}},|x_{B_{i}}|_{1}\in[1..k-1])$ or $(z_{B_{i}}=0^{n^{\prime}},|x_{B_{i}}|_{1}\in[n^{\prime}-k+1..n^{\prime}-1])$. Hence, with the bit values on other blocks fixed, there are at most two solutions in $P$ that are not weakly dominated. Then $m/2$ blocks results in at most $2^{m/2}$ non-weakly-dominated solutions, that is, $|P|\leq 2^{m/2}<3^{m/2}\leq M$. ∎
For ease of reading of the proof sketches in the following sections, we call a Pareto optimum $x$ an inner Pareto optimum if for all blocks $B_{i}=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$, we have $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$. We call $u=(u_{1},n^{\prime}+2k-u_{1},\dots,u_{m/2},n^{\prime}+2k-u_{m/2})$ and $v=(v_{1},n^{\prime}+2k-v_{1},\dots,v_{m/2},n^{\prime}+2k-v_{m/2})$ neighbors if there exists $i\in[1..m/2]$ such that $u_{i}-v_{i}=1$ and $u_{j}=v_{j}$ for all $j\neq i$.
4 SMS-EMOA Can Optimize $m$OJZJ
As discussed before, the original NSGA-II with non-dominated sorting and crowding distance needs at least exponential runtime to cover the full Pareto front of the $m$OneMinMax [ZD23b], which is a special case of $m$OJZJ. Motivated by the question whether there is a simple efficient variant of NSGA-II for many objectives, this section will analyze the runtime of the SMS-EMOA, which still uses the non-dominated sorting but replaces the crowding distance by hypervolume measure.
4.1 Algorithm Description
The SMS-EMOA is a steady-state variant (that is, the offspring size is smaller than the parent size) of the NSGA-II. It maintains a fixed population size $\mu$. For each iteration, only one offspring $x^{\prime}$ is generated from the parent population $P_{t}$. From the $\mu+1$ individuals in the combined population $R_{t}=P_{t}\cup\{x^{\prime}\}$, one individual will be removed. It first sorts the combined population into several fronts $F_{1},\dots,F_{i^{*}}$, where $F_{i}$ contains all mutually non-dominated individuals in $R_{t}\setminus(\bigcup_{j=1}^{i-1}F_{j})$. This is called non-dominated sorting. Different from the original NSGA-II that uses the crowding distance as the secondary criterion to remove the individual, SMS-EMOA removes the one with the smallest hypervolume contribution in the critical front, that is, $F_{i^{*}}$. The hypervolume of the set $S$ w.r.t. a reference point $r$ is defined as
$\operatorname{HV}_{r}(S)=\mathcal{L}\mathopen{}\mathclose{{}\left(\,\bigcup_{u\in S}\{h\in\mathbb{R}^{m}\mid r\leq h\leq f(u)\}}\right),$
where $\mathcal{L}$ is the Lebesgue measure. The hypervolume contribution of an individual $x\in F_{i^{*}}$ is calculated by
$\Delta_{r}(x,F_{i^{*}}):=\operatorname{HV}_{r}(F_{i^{*}})-\operatorname{HV}_{r}(F_{i^{*}}\setminus\{x\}).$
Algorithm 1 shows the general procedure of the SMS-EMOA.
1:Initialize $P_{0}$ by generating $\mu$ solutions uniformly at random from $\{0,1\}^{n}$
2:for $t=0,1,2,\dots,$ do
3:   Select a solution $x$ uniformly at random from $P_{t}$
4:   Generate $x^{\prime}$ by flipping each bit of $x$ independently with probability $1/n$
5:   %Original Population Update
6:   Use fast-non-dominated-sort() [DPAM02] to divide $R_{t}=P_{t}\cup\{x^{\prime}\}$ into $F_{1},\dots,F_{i^{*}}$
7:   Calculate $\Delta_{r}(z,F_{i^{*}})$ for all $z\in F_{i^{*}}$ and find $D=\arg\min_{z\in F_{i^{*}}}{\Delta_{r}(z,F_{i^{*}})}$
8:   Uniformly at random pick $z^{\prime}\in D$ and $P_{t+1}=R_{t}\setminus\{z^{\prime}\}$
9:end for
Algorithm 1 SMS-EMOA
4.2 Runtime of SMS-EMOA
We now establish the runtime of the SMS-EMOA to cover the full Pareto front of the $m$OJZJ. We first give the following lemma to show that for a large enough population size, the function value of any point with rank one in the combined population will survive to the next population. The key of the proof is that all $F_{1}$ points will survive if there is a point with a larger rank, and that at least two points have the same function value if all points are in $F_{1}$ and the population size is large enough.
Lemma 4.
Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$. Specifically, any Pareto front point will be maintained in all future populations once it is reached.
Proof.
Let $x\in F_{1}$. If $\cup_{i\geq 2}F_{i}\neq\emptyset$, then the removal will happen in $\cup_{i\geq 2}F_{i}$, and $x\in P^{t+1}$.
If all solutions are in $F_{1}$, from Lemma 3, we know that $|f(F_{1})|\leq M$. Hence, there exist $z_{1},z_{2}\in F_{1}$ such that $f(z_{1})=f(z_{2})$. Let $S_{x}=\{z\in R_{t}\mid f(x)=f(z)\}$. Then $|S_{x}|\geq 1$. If $|S_{x}|=1$, then $\Delta(x,F_{1})>0=\Delta(z_{1},F_{1})$, and thus $x\in P_{t+1}$. If $|S_{x}|>1$, since the survival selection only removes one solution, then there exists $y\in S_{x}$, such that $y\in P_{t+1}$.
Since any Pareto optimum is always in $F_{1}$, the latter part of this lemma holds. ∎
Recall that the inner Pareto optimum is a Pareto optimum with $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$ for all blocks $B_{i}=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$, as defined before. Since not all points are inner Pareto optimum, the following lemma establishes the runtime to reach one inner Pareto optimum for the first time. The key of the proof is that at most $k$ specific bit flipping of a point that is not an inner Pareto optimum will result in an inner Pareto optimum.
Lemma 5.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $\frac{1}{2}em^{k+1}k^{k}$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
Proof.
This lemma is already proved if the initial population has at least one inner Pareto optimum. Now we consider that the initial population contains no inner Pareto optimum. Consider any population $P$ without inner Pareto optimum. Then we know that there exists a block $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$ such that for all $x\in P$, $A_{i}:=|x_{B_{i}}|_{1}\in[0..k-1]\cup[n^{\prime}-k+1..n^{\prime}]$. We regard it a success if $y$ with $|y_{B_{i}}|_{1}\in[k..n^{\prime}-k]$ is generated. Then the probability to generate such $y$ from $x$ with $A_{i}<k$ is at least
$\binom{n^{\prime}-A_{i}}{k-A_{i}}$ ${}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(k-A_{i})}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k-A_{i}}\geq\frac{1}{e}\binom{n^{\prime}-A_{i}}{k-A_{i}}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k-A_{i}}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}-A_{i}}{n(k-A_{i})}}\right)^{k-A_{i}}$
$\geq{}\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}-k}{nk}}\right)^{k}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}}{2nk}}\right)^{k}=\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{1}{mk}}\right)^{k},$
where the last inequality uses $k\leq n^{\prime}/2$ and the last equality uses $n^{\prime}=2m/n$. We also know that the probability to generate such $y$ from $x$ with $A_{i}>n^{\prime}-k$ is at least
$\binom{A_{i}}{A_{i}-(n^{\prime}-k)}$ ${}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(A_{i}-(n^{\prime}-k))}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{A_{i}-(n^{\prime}-k)}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{A_{i}}{n(A_{i}-(n^{\prime}-k))}}\right)^{A_{i}-(n^{\prime}-k)}$
$\geq{}\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}-k}{nk}}\right)^{k}\geq\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{n^{\prime}}{2nk}}\right)^{k}=\frac{1}{e}\mathopen{}\mathclose{{}\left(\frac{1}{mk}}\right)^{k},$
where the last two inequalities use $n^{\prime}-k<A_{i}\leq n^{\prime}$ and $k\leq n^{\prime}/2$ respectively, and the last equality uses $n^{\prime}=2m/n$. Note that $A_{i}\in[0..k-1]\cup[n^{\prime}-k+1..n^{\prime}]$ holds for all $x\in P$ before such $y$ is generated. Then we know that $y$ will be generated in expected $e(mk)^{k}$ iterations. Since such $y$ cannot be dominated by any solution in $P$, we know that $y\in F_{1}$. From Lemma 4, we know that once $y$ is generated, $f(y)$ will be covered by the next population and by all future populations, and thus future populations will contain at least one solution $z$ with $|z_{B_{i}}|_{1}\in[k..n^{\prime}-k]$.
Since an inner Pareto optimum is reached when $y$ with $|y_{B_{1}}|_{1},\dots,|y_{B_{m/2}}|_{1}\in[k..n^{\prime}-k]$ is generated, we know that the expected iterations to reach an inner Pareto optimum is at most $\frac{m}{2}e(mk)^{k}=\frac{1}{2}em^{k+1}k^{k}$. ∎
We consider the next stage of covering all inner Pareto front points. The following lemma shows the runtime. The key of the proof is that one uncovered Pareto front point can be easily generated from a neighbor Pareto front if it is covered.
Lemma 6.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Assume that the current population contains at least one inner Pareto optimum. Then after at most $en\mu M$ iterations in expectation, all inner Pareto front points are covered.
Proof.
Let $u\in F^{*}$ be an uncovered inner Pareto front point that has some neighbor covered by the population $P$. Since such a neighbor is also a Pareto front point, we know that it is maintained in all future generations. Since the probability to generate an offspring with its function value of $u$ is at least $\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n\mu}\geq\frac{1}{en\mu}$, we know that the expected iterations to witness such an offspring is at most $n\mu$ in expectation. From Lemma 4, we know that once such an offspring is generated, $u$ will be covered in all future populations. Noting that there are at most $M$ inner Pareto front points, we know that in expected at most $en\mu M$ iterations, all Pareto front points are covered. ∎
The last stage is to cover the remaining Pareto front points. The following lemma shows the runtime. The key of the proof is that we can divide the Pareto front into several levels and any individual in the current $i$-th level can be easily generated from a point in the $(i-1)$-th level.
Lemma 7.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Assume that the current population covers all inner Pareto front points. Then after at most $eM\mu n^{k}$ iterations in expectation, the full Pareto front is covered.
Proof.
We call $(k,n^{\prime}+k)$ or $(n^{\prime}+k,k)$ extreme block. We divide the Pareto front points into $m/2+1$ levels, where each point in $i$-th level contains $i$ extreme blocks. Obviously, all inner Pareto front points belong to the 00-th level. Note that any point in $j$-th level ($j\geq 1$) can be generated by a certain point in $(j-1)$-th level via flipping specific $k$ bits with other bits unchanged. Hence, the probability to generate any point in $j$-th level condition on that all $(j-1)$-th level points are covered is at least
$\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-k}\mathopen{}\mathclose{{}\left(\frac{1}{n}}\right)^{k}\geq\frac{1}{e\mu n^{k}}.$
From Lemma 4, we know that such a Pareto front point is reached in at most $e\mu n^{k}$ in expectation, and will survive in all future populations. Consider the process of covering all non-inner Pareto front points in the order of covering the $1$-st level, …, $(m/2)$-th level. Since the 00-th level points are already covered from our assumption, and since there are at most $M$ non-inner Pareto front points, we know that the remaining levels are covered in at most $eM\mu n^{k}$ iterations in expectation. ∎
Summing up the runtime of all stages from Lemmas 5 to 7, we have the following theorem for the runtime of the full coverage of the Pareto front.
Theorem 8.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $\frac{1}{2}em^{k+1}k^{k}+n\mu M+eM\mu n^{k}$ iterations ($\mu+\frac{1}{2}em^{k+1}k^{k}+n\mu M+eM\mu n^{k}$ function evaluations) in expectation, the full Pareto front is covered.
Comparing with the NSGA-II’s exponential lower bound to cover the special case of $m$OJZJ, Theorem 8 shows the efficiency of the SMS-EMOA for many objectives, and indicates that for the NSGA-II, the non-dominated sorting is cool, but the crowding distance as the secondary criterion is deficient.
4.3 Runtime of GSMEO
For an overall comparison, we also show the runtime of the GSMEO on the $m$OJZJ. The GSEMO is a multiobjective counterpart of the $(1+1)$ EA. The initial population contains a uniformly and randomly generated solution. Then in each iteration, a solution is uniformly at random picked from the population to generate an offspring with standard bit-wise mutation. If this offspring cannot be dominated by any solution in the population, it will enter into the population and removes all solutions weakly dominated by it. It is not difficult to see all solutions in the population of the GSEMO are mutually non-dominated. Hence, the population size of the GSEMO is at most $M$ from Lemma 3.
The proof ideas are similar to Lemmas 5 to 7. The only difference is that SMS-EMOA needs Lemma 4 to ensure that all function values covered in $F_{1}$ will remain covered by the next population, but for GSEMO, all such function values are surely covered by the next population as $F_{1}$ contains all mutually non-dominated individuals whose function values will not be removed in the survival selection. Together with the upper bound $M$ of the population size, we have the following theorem.
Theorem 9.
Let $k\leq n^{\prime}/2$. Consider using GSEMO to optimize the $m$OJZJ problem. Then after at most $\frac{1}{2}em^{k+1}k^{k}+nM^{2}+eM^{2}n^{k}$ iterations in expectation, the full Pareto front is covered.
5 Reduced Impact of Stochastic Population Updates
Bian et al. [BZLQ23] proposed a stochastic population update mechanism for the SMS-EMOA and proved that, somewhat unexpectedly given the state of the art, it can lead to significant performance gains. More precisely, they proved that the classic SMS-EMOA with reasonable population size $\mu=\Theta(n)$ solves the OJZJ problem in an expected number of $O(n^{k+2})\cap\Omega(n^{k+1})$ function evaluations, whereas for the SMS-EMOA with stochastic population update, $O(n^{k+2}\min\{1,2^{-k/4}n\})$ suffice. Hence, for $k=\omega(\log n)$, a super-polynomial speed-up is shown.
In this section, we extend the analysis of [BZLQ23] to $m$ objectives. Unfortunately, we will observe that the larger population sizes necessary here reduce the impact of the stochastic population update. As in [BZLQ23], we have no proven lower bounds for the SMS-EMOA with stochastic population update, but our upper-bound proofs suggest that the reduced effect of the stochastic population update on the runtime guarantee is real, that is, the impact on the true runtime is diminishing with the larger population sizes necessary in the many-objective setting.
5.1 Stochastic Population Update
The rough idea of the stochastic selection proposed by Bian et al. [BZLQ23] is that a random half of the individuals survive into the next generation regardless of their quality. The individual to be discarded is chosen from the remaining individuals according to non-dominated sorting and hypervolume contribution as in the classic SMS-EMOA. This approach resembles the random mixing of two acceptance operators in the Move Acceptance Hyper-Heuristic studied recently [LOW23, DDLS23].
More precisely, after generating the offspring $x^{\prime}$, the SMS-EMOA with stochastic selection chooses from $P_{t}\cup\{x^{\prime}\}$ a set $R^{\prime}$ of $\lfloor(\mu+1)/2\rfloor$ solutions randomly with replacement. The individual $z^{\prime}$ to be removed from $P_{t}\cup\{x^{\prime}\}$ is then determined via non-dominated sorting and hypervolume contribution applied to $R^{\prime}$ only. We note that with this mechanism, even the worst solution can enter the next generation with probability at least $1/2$.
5.2 Runtime
To establish the runtime result of the SMS-EMOA with stochastic population update, we separate the key facts into several lemmas for a clear demonstration. The following two lemmas 10 and 11 respectively show the general survival probability of each individual in the combined population, and the survival of the individual with rank one.
Lemma 10.
Consider using SMS-EMOA with stochastic population update to optimize the $m$OJZJ problem. For any $x\in R_{t}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$ with probability at least $1/2$.
Proof.
We simply calculate the probability of $x$ being removed after the population update. Since $x$ is possible to be removed only if it is selected in $R^{\prime}$, we know that it will be removed with probability at most $\lfloor|R^{\prime}|/2\rfloor/|R^{\prime}|\leq 1/2$, and hence $f(x)$ will be maintained in $P_{t+1}$ with probability at least $1/2$. ∎
Lemma 11.
Consider using SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$. Specifically, any Pareto front point will be maintained in all future populations once it is reached.
Proof.
Let $x\in F_{1}$. If $\cup_{i\geq 2}F_{i}\neq\emptyset$, then the removal will happen in $\cup_{i\geq 2}F_{i}$, and $x\in P^{t+1}$.
If all solutions are in $F_{1}$, from Lemma 3, we know that $|f(F_{1})|\leq M$. Since there are $|R^{\prime}|\geq\lfloor(\mu+1)/2\rfloor\geq M+1>M$, we know that there exist $z_{1},z_{2}\in F_{1}$ such that $f(z_{1})=f(z_{2})$. Let $S_{x}=\{z\in R^{\prime}_{t}\mid f(x)=f(z)\}$. Then $|S_{x}|\geq 1$. If $|S_{x}|=1$, then $\Delta(x,F_{1})>0=\Delta(z_{1},F_{1})$, and thus $x\in P_{t+1}$. If $|S_{x}|>1$, since the survival selection only removes one solution, then there exists $y\in S_{x}$, such that $y\in P_{t+1}$.
Since any Pareto optimum is always in $F_{1}$, the latter part of this lemma holds. ∎
Recalling the proof in Lemmas 5 and 6, we know that these lemmas hold merely relies on that $f(F_{1})$ value will be maintained once it is reached. Hence, from Lemma 11, we have the following results for obtaining a population containing at least one inner Pareto optimum, and for covering all inner Pareto front points starting from a population with at least one inner Pareto optimum.
Lemma 12.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. Then
  • after at most $\frac{1}{2}em^{k+1}k^{k}$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum;
  • after at most $en\mu M$ iterations in expectation, all inner Pareto front points are covered starting from a population with at least one inner Pareto optimum.
Now we consider the runtime for the full coverage of the Pareto front after all inner Pareto front points are covered. Similar to the proof of Lemma 7 for the original SMS-EMOA, we partition the Pareto front points into $m/2+1$ levels based on the number of extreme blocks inside. The difference here is that we will consider the effect of the gap points, similar to [BZLQ23].
Lemma 13.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with stochastic population update and with $\mu\geq 2(M+1)$ to optimize the $m$OJZJ problem. Assume that the current population covers all inner Pareto front points. Then after at most $O(\min\{\frac{\mu k^{1/2}}{2^{k/2}},1\}M\mu n^{k})$ iterations in expectation, the full Pareto front is covered.
Proof.
We use the same definitions of the $j$-th ($j\in[0..m/2]$) level where each Pareto front point contains $j$ extreme blocks, and note that 00-th level points are covered. Let $P_{t}$ be the current population. We consider generating any point in $j$-th level condition on that all $(j-1)$-th level points are covered. Let $v$ be an uncovered Pareto front point in the $j$-th level. Then there exists a block $B$ that $v$ has the value of $(k,n^{\prime}+k)$ or $(n^{\prime}+k,k)$. W.l.o.g., let the block value of $v$ is $v_{B}=(k,n^{\prime}+k)$. Besides, there exists a set of points $U$ in the $(j-1)$-th level such that each $u\in U$ is the same as $v$ except that $u$ has the value of $(a,n^{\prime}+2k-a)$ with $a\in[2k..n^{\prime}]$ in the above block $B$. Let $S_{v}=\{f(z)\mid z\in P,z\text{\leavevmode\nobreak\ is the same as\leavevmode\nobreak\ }v\text{\leavevmode\nobreak\ except\leavevmode\nobreak\ }z_{B}\neq(k,n^{\prime}+k)\}$. Obviously, $y:=\min_{z\in S_{v}}z_{B}\leq 2k$. Let $A_{1}$ denote the event $k<y\leq\frac{3}{2}k$, $A_{2}$ be the event $y\geq\frac{7}{4}k$, and $A_{3}$ be $\frac{3}{2}k<y<\frac{7}{4}k$. Consider the potential function
$h(P_{t})=\begin{cases}0,&{}\text{if $v\in P_{t}$},\\ e\mu n^{k/2},&{}\text{if $A_{1}$},\\ e\mu n^{k/2}+1,&{}\text{if $A_{2}\cup A_{3}$}.\end{cases}$
For $A_{1}$, we pessimistically consider the decrease of $h$ by generating $v$ and the increase of $h$ by kicking out a corresponding solution of $y$. Hence, we have
$E[h$ ${}(P_{t})-h(P_{t+1})\mid P_{t},A_{1}]\geq\frac{1}{\mu}\frac{1}{n^{y-k}}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(y-k)}e\mu n^{k/2}-\frac{1}{2}\cdot 1\geq 1-\frac{1}{2}=\frac{1}{2}.$
For $A_{2}$, $h$ will not increase, and there are $y-k\in[\frac{3}{4}k,k]$ ones in the block of a corresponding solution of $y$. We pessimistically consider the decrease of $h$ by flipping $y-k-\frac{1}{2}k$ ones and keeping other bits unchanged of the corresponding solution. Hence, we have
$E[h$ ${}(P_{t})-h(P_{t+1})\mid P_{t},A_{2}]\geq\frac{1}{\mu}\frac{\binom{y-k}{y-k-k/2}}{n^{y-k-k/2}}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-(y-k-k/2)}\frac{1}{2}$
[…]
where the last inequality uses $\frac{3}{4}k\leq y-k\leq k$. For $A_{3}$, $h$ will not increase, and there are $y-k\in(\frac{1}{2},\frac{3}{4}k)$ ones in the block of a corresponding solution of $y$. We pessimistically consider the decrease of $h$ by flipping $\frac{1}{4}k$ ones and keeping other bits unchanged of the corresponding solution. Hence, we have
$E[h$ ${}(P_{t})-h(P_{t+1})\mid P_{t},A_{3}]\geq\frac{1}{\mu}\frac{\binom{y-k}{k/4}}{n^{k/4}}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-k/4}\frac{1}{2}\geq\frac{1}{2e\mu}\frac{\binom{k/2}{k/4}}{n^{k/4}}.$
Therefore, we have
$E[h$ ${}(P_{t+1})-h(P_{t})\mid P_{t},v\notin P_{t}]=\sum_{i=1}^{3}\Pr[A_{i}]E[h(P_{t+1})-h(P_{t})\mid P_{t},A_{i}]$
[…]
where the third inequality uses $\binom{k/2}{k/4}\leq\binom{\frac{3}{4}k}{k/2}$ and $n^{k/2}\geq n^{k/4}$, and the last inequality uses $k\geq 1$. Since the maximum $h$ is $e\mu n^{k/2}+1$, we know that the Pareto front point $v$ will be reached in at most
$(e\mu n^{k/2}+1)\frac{{4e\mu n^{k/2}\sqrt{k}}}{4^{k/4}}=O\mathopen{}\mathclose{{}\left(\frac{\mu k^{1/2}}{2^{k/2}}\mu n^{k}}\right)$
in expectation, and will survive in all future populations. Consider the process of covering all non-inner Pareto front points in the order of covering the $1$-st level, …, $(m/2)$-th level. Since the 00-th level points are already covered from our assumption, and since there are at most $M$ non-inner Pareto front points, we know that the remaining levels are covered in at most $O(\frac{\mu k^{1/2}}{2^{k/2}}M\mu n^{k})$ iterations in expectation. With the same argument in the proof of Lemma 7, we know that the expected number of iterations to cover the full Pareto front is at most $eM\mu n^{k}$. Hence, we proved this lemma. ∎
Combining Lemmas 12 and 13, we have the runtime of the SMS-EMOA with stochastic population update in the following theorem.
Theorem 14.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq 2(M+1)$ and stochastic population update to optimize the $m$OJZJ problem. Then after at most $\frac{1}{2}em^{k+1}k^{k}+en\mu M+O(\min\{\frac{\mu k^{1/2}}{2^{k/2}},1\}M\mu n^{k})$ iterations in expectation, the full Pareto front is covered.
Theorem 14 shows that the best runtime for the SMS-EMOA with stochastic population update is $O(\min\{\frac{Mk^{1/2}}{2^{k/2}},1\}M^{2}n^{k})$ when we choose $\mu=\Theta(M)$. Recall that the runtime of the original SMS-EMOA is $O(M^{2}n^{k})$ for $\mu=\Theta(M)$ from Theorem 8. We know the asymptotic speed-up is $\Theta(\max\{\frac{2^{k/2}}{Mk^{1/2}},1\})$ with $M=(2n/m-2k+3)^{m/2}$. Hence, when the objective number is not small, like at least the same size of $k$, then the speed-up is only $\Theta(1)$. We note that it is not contradictory to the nice speed-up obtained for two objectives in [BZLQ23] but adds a limitation situation for the stochastic population update.
6 Heavy-Tailed Mutation Helps
In the previous section, we saw that the advantages of a stochastic population update do not well generalize from bi-objective optimization to many-objective optimization. We now regard another design choice that so far was only analyzed in bi-objective optimization, namely a heavy-tailed mutation operator. We shall prove that the $k^{\Omega(k)}$ factor speed-up observed in bi-objective optimization extends to many objectives.
6.1 Heavy-Tailed Mutation
Different from the standard bit-wise mutation, the heavy-tailed mutation flips each bit of the solution with probability $\alpha/n$ independently, where $\alpha$ follows a power-law distribution with parameter $\beta$ defined as follows.
Definition 15.
Let $n\in\mathbb{N}$ and $\beta>1$. We say $\alpha$ follows the power-law distribution if for all $i\in[1..n/2]$, $\Pr[\alpha=i]=\mathopen{}\mathclose{{}\left(\sum_{j=1}^{n/2}k^{-\beta}}\right)^{-1}i^{-\beta}$.
This power-law distribution has a good feature that a point with Hamming distance of $j$ of a given bit-string will be generated with a probability of $\Omega(j^{-\beta})$. It will facilitate the jump with a larger size and facilitate escaping from the local optimum. The heavy-tailed mutation has resulted in an asymptotic performance speed-up by a factor of $k^{\Omega(k)}$ of the $(1+1)$ EA on the single-objective Jump function with gap size $k$ [DLMN17]. For the bi-objective OJZJ, the speed-up of $k^{\Omega(k)}$ is also witnessed when it is applied to the GSEMO [DZ21] and to the NSGA-II [DQ23a]. Several other positive theoretical results exist for this heavy-tailed mutation, or more generally, other heavy-tailed parameter choises [FQW18, WQT18, QGWF21, COY21, DELQ22, ABD22, DGI22, DR23]. Prior to this work, no theoretical analysis of the heavy-tailed mutation operator in many-objective optimization existed.
6.2 Runtime
We now analyze the runtime of the SMS-EMOA using the heavy-tailed mutation operator instead of bit-wise mutation, the standard choice for this algorithm.
Since the heavy-tailed mutation does not change the survival selection of the original SMS-EMOA, we easily have the following corollary about the survival of the individuals with rank one from Lemma 4.
Corollary 16.
Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem. For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$. Specifically, any Pareto front point will be maintained in all future populations once it is reached.
With similar proof ideas in Lemma 5 to 7 but noticing the probability of generating a solution with a specific Hamming distance, we obtain the following lemmas for reaching the first inner Pareto optimum (Lemma 17), covering all inner Pareto front points (Lemma 18), and covering the remaining Pareto front points (Lemma 19).
Lemma 17.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $C_{n/2}^{\beta}\binom{n}{k}mk^{\beta}$ iterations in expectation, the population (and also the populations afterwards) contains at least one inner Pareto optimum.
Proof.
We use similar arguments of Lemma 5 except following probability calculations. The probability to generate a $y$ with $|y_{B_{i}}|_{1}\in[k..n^{\prime}-k]$ from $x$ with $A_{i}<k$ or with $A_{i}>n^{\prime}-k$ is at least $\binom{n}{k}^{-1}P_{k}^{\beta}\geq\binom{n}{k}^{-1}\mathopen{}\mathclose{{}\left(C_{n/2}^{\beta}}\right)^{-1}k^{-\beta}$. Hence such $y$ is generated in expected $\binom{n}{k}C_{n/2}^{\beta}k^{\beta}$ iterations. Hence the expected iterations to reach an inner Pareto optimum is at most $\frac{m}{2}\binom{n}{k}C_{n/2}^{\beta}k^{\beta}=\binom{n}{k}C_{n/2}^{\beta}mk^{\beta}$ iterations. ∎
Lemma 18.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem. Assume that the current population contains at least one inner Pareto optimum. Then after at most $\frac{e\beta}{\beta-1}n\mu M$ iterations in expectation, all inner Pareto front points are covered.
Proof.
We use similar arguments of Lemma 6 except following probability calculation. The probability to generate an offspring with its function value of an uncovered inner Pareto front point from a neighbor covered by the population is at least $\frac{1}{n\mu}P_{1}^{\beta}=\frac{\beta-1}{e\beta n\mu}$. ∎
Lemma 19.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with heavy-tailed mutation and with $\mu\geq M$ to optimize the $m$OJZJ problem. Assume that the current population covers all inner Pareto front points. Then after at most $\binom{n}{k}C_{n/2}^{\beta}\mu Mk^{\beta}$ iterations in expectation, the full Pareto front is covered.
Proof.
We use similar arguments of Lemma 7 except following probability calculation. The probability to generate any point in $j$-th level condition on that all $(j-1)$-th level points are covered is at least
$\frac{1}{\mu}\binom{n}{k}^{-1}P_{k}^{\beta}\geq\binom{n}{k}^{-1}\mathopen{}\mathclose{{}\left(C_{n/2}^{\beta}}\right)^{-1}\frac{1}{\mu}k^{-\beta}.\qed$
Considering the whole process for the full coverage of the Pareto front from Lemmas 17 to 19, we have the following runtime result for the SMS-EMOA with heavy-tailed mutation.
Theorem 20.
Let $k\leq n^{\prime}/2$. Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. Then after at most $O(M\mu k^{\beta-0.5-k}n^{k})$ iterations or function evaluations in expectation, the full Pareto front is covered.
Theorem 20 shows the best asymptotic upper bound of the runtime of the SMS-EMOA with heavy-tailed mutation is $O(M^{2}n^{k}k^{\beta-0.5-k})$. Comparing with the runtime of $O(M^{2}n^{k})$ for $\mu=\Theta(M)$ for the original SMS-EMOA from Theorem 8, we see a speed-up of $k^{k+0.5-\beta}$, which is also the speed-up reached for the single objective and two objectives.
7 SMS-EMOA Also Performs Well for Bi-objective Optimization
The above sections discussed the runtime of the SMS-EMOA for many objectives, and showed that it performs well for many objectives, which is different from the original NSGA-II. Since the only theory paper [BZLQ23] merely considers its performance for the bi-objective OJZJ, in this section, we will analyze its runtime for other well-known bi-objective benchmarks to have a more comprehensive theoretical understanding.
7.1 OneMinMax and LOTZ
OneMinMax [GL10] (also the CountingOnesCountingZeroes [${}^{+}$02]) and LOTZ [LTZ04] (also the weighted version WLPTNO [QYZ13]) are the bi-objective counterpart of the well-analyzed single-objective OneMax and LeadingOnes respectively. For OneMinMax, two objectives count the number of ones and zeros in a given bit-string respectively. For LOTZ, the first objective counts the continuous ones starting from the first bit position, and the second one counts the continuous zeros starting from the last bit position. Both benchmarks with a simple and clear structure facilitate the fundamental theoretical understanding of the MOEAs. We note that there are other benchmarks [QTZ16, LZZZ16, DOSS23b], but here we focus on the two kinds of most popular benchmarks for analysis.
Here are formal definitions of the OneMinMax and LOTZ.
Definition 21 ([GL10]).
Let $n\in\mathbb{N}$ be the problem size. The OneMinMax function :$:\{0,1\}^{n}\rightarrow\mathbb{R}$ is defined by
$\textsc{OneMinMax}(x)=\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}(1-x_{i}),\sum_{i=1}^{n}x_{i}}\right)$
for any $x=(x_{1},\dots,x_{n})\in\{0,1\}^{n}$.
Definition 22 ([LTZ04]).
Let $n\in\mathbb{N}$ be the problem size. The LOTZ function :$:\{0,1\}^{n}\rightarrow\mathbb{R}$ is defined by
$\textsc{LOTZ}(x)=\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}\prod_{j=1}^{i}x_{j},\sum_{i=n}^{1}\prod_{j=n}^{i}x_{j}}\right)$
for any $x=(x_{1},\dots,x_{n})\in\{0,1\}^{n}$.
7.2 Runtime
We now establish the runtime of the SMS-EMOA on the OneMinMax and LOTZ. Note that the Pareto front sizes for OneMinMax and LOTZ are $n+1$. Hence, with the similar proof of Lemma 4, we know that for $\mu\geq n+1$ and for a specific function value $u$ reached by $F_{1}$, one individual in the non-empty $\cup_{i\geq 2}F_{i}$ or one individual $x$ in $F_{1}$ with $\Delta(x,F_{1})=0$ will be removed and $u$ will remain covered in the next generation. We have the following formal description.
Lemma 23.
Consider using SMS-EMOA with $\mu\geq n+1$ to optimize the LOTZ or OneMinMax problem. For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$. Specifically, any Pareto front point will be maintained in all future populations once it is reached.
Theorem 24 shows the runtime of SMS-EMOA on OneMinMax. The proof considers the probability to reach an uncovered Pareto front point from a specific covered one. From Lemma 23 that any covered Pareto front point will be kept, we then turn such probability to the runtime argument.
Theorem 24.
Consider using SMS-EMOA with $\mu\geq n+1$ to optimize the OneMinMax problem with problem size $n$. Then after at most $2e\mu n(\ln n+1)$ iterations in expectation, the full Pareto front is covered.
Proof.
Since each search point in $\{0,1\}^{n}$ is Pareto optimal, from Lemma 23, we know that any Pareto front point covered by the population $P_{t}$ will be maintained in all future generations. Assume that $(i,n-i)$ for some $i\in[0..n]$ is covered in $P_{t}$. Then the probability to reach a solution with function value $(i+1,n-i-1)$ is at least
$\frac{1}{\mu}\frac{n-i}{n}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\geq\frac{n-i}{e\mu n},$
and the probability to reach a solution with function value $(i-1,n-i+1)$ is at least
$\frac{1}{\mu}\frac{i}{n}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\geq\frac{i}{e\mu n}.$
Thus the expected numbers of iterations to cover $(i+1,n-i-1)$ and $(i-1,n-i+1)$ are at most $\frac{e\mu n}{n-i}$ and at most $\frac{e\mu n}{i}$ respectively. Hence, the expected number of iterations to cover the full Pareto front is
$\Bigg{(}\sum_{j=i}^{1}\frac{e\mu n}{n-j}\Bigg{)}+\mathopen{}\mathclose{{}\left(\sum_{j=i}^{n-1}\frac{e\mu n}{j}}\right)\leq 2\sum_{j=1}^{n-1}\frac{e\mu n}{j}\leq 2e\mu n(\ln n+1).\qed$
We now establish the runtime of SMS-EMOA on LOTZ, see Theorem 25. In the proof, the process is divided into two stages. The first stage is to cover a Pareto front point for the first time and the second stage is to cover the full Pareto front starting from one covered Pareto front point. For the first stage, we pessimistically consider the time to reach $1^{n}$ and for the second stage, we consider the time to reach $01^{i}0^{n-i},i=n-1,\dots,$ 0 one by one.
Theorem 25.
Consider using SMS-EMOA with $\mu\geq n+1$ to optimize the LOTZ problem with problem size $n$. Then after at most $2e\mu n^{2}$ iterations in expectation, the full Pareto front is covered.
Proof.
We first consider the time to reach $1^{n}$ in the first stage. Let $S=\arg\max_{x\in P_{t}}f_{1}(x)$ and $S^{\prime}=\arg\max_{x\in R_{t}}f_{1}(x)$. Then there is at least one individual $x\in S^{\prime}$ such that $x\in F_{1}$ (If not, there is $y\notin S^{\prime}$, such that $x\prec y$ for all $x\in S^{\prime}$. That is, $f_{1}(y)\geq f_{1}(x),f_{2}(y)\geq f_{2}(x),$ and at least one of the inequality is strict. Since $y\notin S$, we know that $f_{1}(y)<f_{1}(x)$, which contradicts $f_{1}(y)\geq f_{1}(x)$). From Lemma 23, we know that $f(x)$ (also $f_{1}(x)$) will be covered by $P_{t+1}$. Hence, to reach $1^{n}$, we only need to reach such $S^{\prime}$ with $f_{1}(S^{\prime})=n$. We calculate
$\Pr[f_{1}(S^{\prime})>f_{1}(S)\mid f_{1}(S)]\geq\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n}\geq\frac{1}{e\mu n}$
where we consider choosing one individual of $S$ and flipping the critical bit value from 00 to $1$. Since $f_{1}(S)$ will not decrease, we know that the expected iterations to witness such $S^{\prime}$ is at most $e\mu n$. Since there are at most $n$ such $S^{\prime}$ with $f_{1}(S^{\prime})>f_{1}(S)$ to reach $1^{n}$, we know that $1^{n}$ will be reached in at most $e\mu n^{2}$ in expectation.
Once $1^{n}$ is reached, we consider the time to reach $1^{n-1}0,\dots,10^{n-1},0^{n}$ one by one. Assume that $1^{i}0^{n-i}$ with $i\in[1..n]$ is reached. Then to reach $1^{i-1}0^{n-i+1}$, we only need to choose $1^{i}0^{n-i}$ from $P_{t}$, and flip the $i$-th bit from $1$ to 00 with other bits unchanged, which happens with probability
$\frac{1}{\mu}\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{n-1}\frac{1}{n}\geq\frac{1}{e\mu n}.$
From Lemma 23, we know that any Pareto front point will be maintained in all future generations once reached. Hence, $1^{i}0^{n-i}\in P_{\tau}$ for all $\tau\geq t$. Thus, the expected number of iterations to reach $1^{i-1}0^{n-i+1}$ is at most $e\mu n$. Since we have $n$ other Pareto optima to reach, we know that the expected number of iterations to cover all remaining Pareto front points is at most $e\mu n^{2}$.
In summary, the full Pareto front will be covered in at most $2e\mu n^{2}$ iterations in expectation. ∎
Note that the runtime of the GSEMO and NSGA-II are $O(n^{2}\ln n)$ and $O(n^{3})$ for the OneMinMax and LOTZ respectively. Hence, Theorems 24 and 25 shows the SMS-EMOA achieves the same asymptotic runtime as GSEMO and NSGA-II.
8 Conclusion
Motivated by the observed difficulty of the NSGA-II for many objectives, this paper resorted to the SMS-EMOA, a variant of the steady-state NSGA-II, and proved that, different from the NSGA-II, it efficiently solves the $m$OJZJ problem. Noting that the SMS-EMOA also employs non-dominated sorting, but replaces the crowding distance with the hypervolume, this result together with the one of Wietheger and Doerr [WD23] supports the conclusion that non-dominated sorting is a good building block for MOEAs, but the crowding distance has deficiencies for more than two objectives.
We also showed that the stochastic population update proposed in [BZLQ23] for the bi-objective SMS-EMOA becomes less effective for many objectives. All these results in a very rigorous manner support the general knowledge that multi-objective optimization becomes increasingly harder with growing numbers of objectives.
On the positive side, we showed that the advantages of heavy-tailed mutation, previously observed in single- and bi-objective optimization, remain when increasing the number of objectives.
Given that our first result shows advantages of the SMS-EMOA and there is only one previous work performing rigorous runtime analyses for this algorithm, we extended our knowledge in this direction by proving competitive runtime guarantees for this algorithm on the two most prominent bi-objective benchmarks.
Acknowledgments
This work was supported by National Natural Science Foundation of China (Grant No. 62306086), Science, Technology and Innovation Commission of Shenzhen Municipality (Grant No. GXWD20220818191018001), Guangdong Basic and Applied Basic Research Foundation (Grant No. 2019A1515110177). This research benefited from the support of the FMJH Program Gaspard Monge for optimization and operations research and their interactions with data science.